In a terrestrial frame of reference, a star is 56 light-years away. How fast would a person have to travel so that the distance to him is only 35 light-years?


1
Expert's answer
2022-03-28T16:15:26-0400

"v=c\sqrt{1-\frac {l^2}{l_0^2}}=0.78c."